3. It can be shown that Use this formula to find 9 +18+27 +36++5562. (5 pts)
Answer #1

Let S = 9 +18 + 27 + 36 +....+ 5562
S = 9(1 + 2 + 3 + 4 + 5 +....+ 618)
1 + 2 + 3 + 4 + 5+...+ 618 = 618(618+D-191271
S = 9 x 191271 = 1721439